Find the equation of the given linear function.
f(x) =
x
−15 0 15 20
f(x)
−2 −5 −8 −9
for your better understanding the numbers were in a table and it was numbers of x then the numbers of f(x)

Answer by Alan3354(69443)

Find the equation of the given linear function.
f(x) =
x
−15 0 15 20
f(x) = y
−2 −5 −8 −9
--------------------------
Assuming it is linear, pick any 2 of the points.
eg, (0,-5) and 15,-8)
Find the slope m:
m = diffy/diffx = -3/15 = -1/5
----
Use y-y1 = m*(x-x1) where (x1,y1) is any of the given points.
=======================
I would check all the points to see if they're on the line - satisfy the equation.